# Efficient Quantification of Time-Series Prediction Error: Optimal Selection Conformal Prediction

### v2

@@ IX-B Detailed Experiment Results and Visualizations
CopulaCPTS [sun2024copula] 73.2 $\pm$ 8.0 93.45 $\pm$ 17.03 77.1 $\pm$ 6.6 104.86 $\pm$ 14.86 81.4 $\pm$ 6.3 116.53 $\pm$ 14.42
CRD-Ell [Tumu2024multi-modal] 85.9 $\pm$ 5.6 1206.44 $\pm$ 238.9 91.2 $\pm$ 4.9 1486.12 $\pm$ 365.38 96.7 $\pm$ 3.3 2074.58 $\pm$ 549.95
CRD-Rect [Tumu2024multi-modal] 85.9 $\pm$ 5.7 798.16 $\pm$ 167.12 91.3 $\pm$ 4.8 999.14 $\pm$ 277.5 77.3 $\pm$ 39.2 nan $\pm$ nan
CF-RNN [stankeviciute2021conformal] 92.8 $\pm$ 3.8 165.4 $\pm$ 10.66 92.8 $\pm$ 3.8 165.4 $\pm$ 10.66 92.8 $\pm$ 3.8 165.4 $\pm$ 10.66
MC-dropout [gal2016dropout] 12.8 $\pm$ 4.7 7.48 $\pm$ 0.61 15.6 $\pm$ 4.9 8.55 $\pm$ 0.7 19.3 $\pm$ 5.0 10.19 $\pm$ 0.83
@@ IX-C Further Discussions on the Experiment Results
For deep UQ methods, MC-dropout [gal2016dropout] does not have validity guarantees and is thus far below the reference target levels for Particle ($\sigma=0.05$), Drone, and Covid-19 datasets; CF-RNN [stankeviciute2021conformal] is one of the classic works of CP in time series which holds the validity guarantee, but it offers overly large confidence regions. For the baselines from newer works, CRD [Tumu2024multi-modal] is unstable and produces excessive large confidence sets that cannot be shown in Figure 4. We hypothesize that this is due to the error accumulation during the fitting and estimation procedures of the CRD method. The performance of CopulaCPTS [sun2024copula] is also unstable. It has large variances for some cases (see the area graph with confidence = 85%, 90% & 95% in Figure 3), and also has unstable coverages (see results on Covid-19 dataset). Among all baselines, LCP [cleaveland2024conformal] has the best performance as it achieves target coverages while having smaller volumes than other baselines. However, the computation cost for LCP can be very large (see Table I in Section IV-C), and it still has larger volumes than our method, OSCP.
@@
A general conclusion from the results in Table II and Figure 3 & 4 is that our method with $l_{2}$-norm, OSCP-$l_{2}$, outperforms all the baselines on 4 case studies for all 10 confidence levels ($0.5\%$ to $0.95\%$). Specifically, OSCP-$l_{2}$ has coverages higher than all the target confidence levels, and it achieves volume reductions (average value over 10 confidence levels) of 16.03%, 14.32%, 14.01%, 16.93% compared to previous SOTA (LCP, [cleaveland2024conformal]) on Particle ($\mathbf{\sigma=0.01}$), Particle ($\mathbf{\sigma=0.05}$), Drone, Covid-19 datasets, respectively.
Besides, our method with ellipsoidal norm, OSCP-$Ell$, achieves volume reduction of 13.65%, 14.81%, 33.5% compared to previous SOTA (LCP, [cleaveland2024conformal]) on Particle ($\mathbf{\sigma=0.01}$), Particle ($\mathbf{\sigma=0.05}$), Drone, respectively (since there is no ellipsoid for 1d case, we did not run OSCP-$Ell$ on the Covid-19 dataset). Ellipsoidal norm fitted an ellipsoid shape for each time step using $D_{\mathrm{cal,1}}$, and ellipsoid norm performs better when the residuals exhibit greater non-Gaussianity and/or when the dimension $d$ is larger. This explains why we can observe a more significant volume reduction on Drone dataset ($d=3$) using OSCP-$Ell$ (33.5%) compared to OSCP-$l_{2}$ (14.1%), while we cannot see significant differences on the Particle datasets (13.65% vs. 16.03% and 14.81% vs. 14.32%, respectively). In a word, when $|D_{\mathrm{cal}}|$ is small and we have no prior knowledge of the behavior of the residual at each time step, we can just assume the residual at each time step is Gaussian-error and use OSCP-$l_{2}$. Otherwise, if we know that the residual is highly non-Gaussian and the $|D_{\mathrm{cal}}|$ is not too small, using OSCP-$Ell$ is more preferable.
@@ IX-D Construction of Ellipsoid-Norm
We wish to highlight that the ellipsoid norm is fitted for each time step individually, so we have $T$ different ellipsoidal norms in total, which ensures a locally-adaptive shape fitting. To guarantee robust construction of ellipsoid-norms against numerical issues, one can use pseudo-inverses of covariance matrices via the singular-value-decomposition technique instead of the standard inverses (see [Xu2024Conformal]).

### v3

@@ IX-B Detailed Experiment Results and Visualizations
CopulaCPTS [22] 73.2 $\pm$ 8.0 93.45 $\pm$ 17.03 77.1 $\pm$ 6.6 104.86 $\pm$ 14.86 81.4 $\pm$ 6.3 116.53 $\pm$ 14.42
CRD-Ell [26] 85.9 $\pm$ 5.6 1206.44 $\pm$ 238.9 91.2 $\pm$ 4.9 1486.12 $\pm$ 365.38 96.7 $\pm$ 3.3 2074.58 $\pm$ 549.95
CRD-Rect [26] 85.9 $\pm$ 5.7 798.16 $\pm$ 167.12 91.3 $\pm$ 4.8 999.14 $\pm$ 277.5 77.3 $\pm$ 39.2 nan $\pm$ nan
CF-RNN [21] 92.8 $\pm$ 3.8 165.4 $\pm$ 10.66 92.8 $\pm$ 3.8 165.4 $\pm$ 10.66 92.8 $\pm$ 3.8 165.4 $\pm$ 10.66
MC-dropout [7] 12.8 $\pm$ 4.7 7.48 $\pm$ 0.61 15.6 $\pm$ 4.9 8.55 $\pm$ 0.7 19.3 $\pm$ 5.0 10.19 $\pm$ 0.83
@@ IX-C Further Discussions on the Experiment Results
For deep UQ methods, MC-dropout [7] does not have validity guarantees and is thus far below the reference target levels for Particle ($\sigma=0.05$), Drone, and Covid-19 datasets; CF-RNN [21] is one of the classic works of CP in time series which holds the validity guarantee, but it offers overly large confidence regions. For the baselines from newer works, CRD [26] is unstable and produces excessive large confidence sets that cannot be shown in Figure 4. We hypothesize that this is due to the error accumulation during the fitting and estimation procedures of the CRD method. The performance of CopulaCPTS [22] is also unstable. It has large variances for some cases (see the area graph with confidence = 85%, 90% & 95% in Figure 3), and also has unstable coverages (see results on Covid-19 dataset). Among all baselines, LCP [5] has the best performance as it achieves target coverages while having smaller volumes than other baselines. However, the computation cost for LCP can be very large (see Table I in Section IV-C), and it still has larger volumes than our method, OSCP.
@@
A general conclusion from the results in Table II and Figure 3 & 4 is that our method with $l_{2}$-norm, OSCP-$l_{2}$, outperforms all the baselines on 4 case studies for all 10 confidence levels ($0.5\%$ to $0.95\%$). Specifically, OSCP-$l_{2}$ has coverages higher than all the target confidence levels, and it achieves volume reductions (average value over 10 confidence levels) of 16.03%, 14.32%, 14.01%, 16.93% compared to previous SOTA (LCP, [5]) on Particle ($\mathbf{\sigma=0.01}$), Particle ($\mathbf{\sigma=0.05}$), Drone, Covid-19 datasets, respectively.
Besides, our method with ellipsoidal norm, OSCP-$Ell$, achieves volume reduction of 13.65%, 14.81%, 33.5% compared to previous SOTA (LCP, [5]) on Particle ($\mathbf{\sigma=0.01}$), Particle ($\mathbf{\sigma=0.05}$), Drone, respectively (since there is no ellipsoid for 1d case, we did not run OSCP-$Ell$ on the Covid-19 dataset). Ellipsoidal norm fitted an ellipsoid shape for each time step using $D_{\mathrm{cal,1}}$, and ellipsoid norm performs better when the residuals exhibit greater non-Gaussianity and/or when the dimension $d$ is larger. This explains why we can observe a more significant volume reduction on Drone dataset ($d=3$) using OSCP-$Ell$ (33.5%) compared to OSCP-$l_{2}$ (14.1%), while we cannot see significant differences on the Particle datasets (13.65% vs. 16.03% and 14.81% vs. 14.32%, respectively). In a word, when $|D_{\mathrm{cal}}|$ is small and we have no prior knowledge of the behavior of the residual at each time step, we can just assume the residual at each time step is Gaussian-error and use OSCP-$l_{2}$. Otherwise, if we know that the residual is highly non-Gaussian and the $|D_{\mathrm{cal}}|$ is not too small, using OSCP-$Ell$ is more preferable.
@@ IX-D Construction of Ellipsoid-Norm
We wish to highlight that the ellipsoid norm is fitted for each time step individually, so we have $T$ different ellipsoidal norms in total, which ensures a locally-adaptive shape fitting. To guarantee robust construction of ellipsoid-norms against numerical issues, one can use pseudo-inverses of covariance matrices via the singular-value-decomposition technique instead of the standard inverses (see [29]).